An investor bought 400 shares of stock some at $6.75 per share and some at $5.00 per share. If the total cost was $2306.25 how many shares of each stock did the investor buy? Round to two decimal places if necessary
Answer by Boreal(15235)

x shares at 6.75 and (400-x) at 5
x*(6.75)+(400-x)*5=2306.25
6.75x+2000-5x=2306.25
1.75x=306.25
x=175 shares @$6.75=$1181.25
(400-x)=225 shares@$5=$1125
Total is $2306.25